# High-throughput polarization-independent spatial mode shaping with mode-selective photonic lanterns

### v1

@@ Abstract
Figure 2: (A). 20 random mode interference profiles obtained at the MM end of the PL for various piezo voltage combinations. The frames are extracted from the on-the-fly video file during the voltage scan. As seen, the mode energy is concentrated in Gaussian-like profiles on various voltage combinations. The dotted circle outlines the approximate core region of the multimode end. Images share the same spatial scale, scale bar, 0.5 mm. (B) Respective mode energies in the image frames, quantified by summing the pixels within the core region. (C) 3D visualization of 6 random images with smoothed surface rendering and corresponding images projected on YZ wall, capturing the diverse spatial evolution of the light field upon accousto-optic scan. The final frame exhibits a Gaussian-like intensity distribution, indicative of effective multimode beam shaping and high-fidelity conversion to a desired single-mode spatial profile.
Figure 3: (A) Representative mode intensity profiles extracted from the recorded video sequence, illustrating key interference outcomes. Subfigure (a) corresponds to the frame exhibiting maximum total photon energy across the MM core. Subfigures (b–d) present distinct instances of Gaussian-like mode interference patterns with varying degrees of spatial localisation. Subfigure (e) highlights a highly non-Gaussian interference state, while (f) shows the mode profile associated with the minimum total energy. Scale bar, 0.5 mm. (B) Total photon counts for each of the six frames, calculated by integrating pixel intensities within the defined MM core region. Corresponding mode profiles are shown as insets for direct comparison between spatial distribution and total energy. The bar plot shows the energy conversion efficiencies of the six mode profiles (subfigures (a–f)), calculated by normalizing their photon fluxes to that of subfigure (f), which exhibits the highest photon flux in the entire video. This comparison illustrates the variation in energy conversion efficiency across different spatial interference patterns.

### v3

@@ Abstract
Figure 2: (A). 20 random mode interference profiles obtained at the MM end of the PL for various piezo voltage combinations. The frames are extracted from the on-the-fly video file during the voltage scan. As seen, the mode energy is concentrated in Gaussian-like profiles on various voltage combinations. The dotted circle outlines the approximate core region of the multimode end. All images cover the same spatial area, scale bar, 0.5 mm. (B) Respective mode energies in the image frames, quantified by summing the pixels within the core region. (C) 3D visualization of 6 random images with smoothed surface rendering and corresponding images projected on YZ wall, capturing the diverse spatial evolution of the light field upon accousto-optic scan. The final frame exhibits a Gaussian-like intensity distribution, indicative of effective multimode beam shaping and high-fidelity conversion to a desired single-mode spatial profile.
Figure 3: (A) Representative mode intensity profiles extracted from the recorded video sequence, illustrating key interference outcomes. Subfigure (a) presents the image frame with the minimum photon flux, subfigure (b) highlights a highly non-Gaussian interference state, subfigures (c-f) present image frames with distinct instances of Gaussian-like mode interference patterns. Subfigure (f)- the mode profile associated with the maximum photon flux. Scale bar, 0.5 mm. (B) Total photon counts for each of the six frames, calculated by integrating pixel intensities within the defined MM core region. Corresponding mode profiles are shown as insets for direct comparison between spatial distribution and total energy. The bar plot shows the energy conversion efficiencies of the six mode profiles (subfigures (a–f)), calculated by normalizing their photon fluxes to that of subfigure (f), which exhibits the highest photon flux in the entire video. This comparison illustrates the variation in energy conversion efficiency across different spatial interference patterns.